40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12108     Entry time: Thu May 5 14:05:01 2016     In reply to: 12105
Author: rana 
Type: Update 
Category: endtable upgrade 
Subject: ALS status update 

All seems very fishy. Its not good to put attenuators and filters in nilly-willy.

  1. Once the post-PD bandpass has been designed and constructed, you should be able to use whatever PD gain setting gives you the best SNR. There's no need to use more PD gain than necessary; it just reduces the PD bandwidth. What is the input referred current noise of the PD at the different gain settings?
  2. The open loop mixer output *should* be very large. It should be reduced to mV only when the loop is closed.
  3. The better way to estimate the modulation depth is to lock the arm on red as usual and then scan the EX laser and look at the green transmission. The FSR is 3.7 MHz, so the SBs should show up well in a narrow scan around the carrier.
  4. I guess its going to be tough to impedance match the splitter box to the NPRO PZT, since its impedance is all over the place at 200-300 kHz, but you could put a 50 Ohm in-line terminator in there somewhere?
  5. The Bode plot seems to indicate that we could easily get a 10 kHz UGF and then switch on a Boost. Is the remote Boost switch disabled or always ON? I am suspicious of the plot and think that the coarse trace is probably missing some sharp resonances which will sneakily bite you.
ELOG V3.1.3-